OpenStudy (anonymous):

Find equations of the tangent lines to the curves y=f(x) andy=g(x) at x=a. Show that these tangent lines intersect on the x-axis at the point (a+ 1=c;0)

OpenStudy (dumbcow):

equation of tangent line is given as: \[y - f(a) = f'(a)(x-a)\] \[\large f'(x) = -c e^{-cx}\] \[\large y = (-c e^{-ac})(x-a) + e^{-ac}\] the same is done for g(x) in same way then set the 2 tangent lines equal to find intersection point

OpenStudy (dumbcow):

\[\large \large (-c e^{-ac})(x-a) + e^{-ac}= (c e^{-ac})(x-a) - e^{-ac}\] which simplifies to \[\large 2c e^{-ac} x =2ace^{-ac}+ 2e^{-ac}\] \[x = a + \frac{1}{c}\]